OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 4 sin^2x - 4 sin x + 1 = 0 I've factored but I don't know what to do from here: Where do I go from (2sin-1)^2 = 0?

OpenStudy (anonymous):

2sinx - 1 = 0 Sinx = 1/2 X = pi/6 Or x = 5Pi/6